# Corks, exotic 4-manifolds and genus functions

### v3

@@ 6.1. Definitions and examples
Suppose a 4-dimensional 1-handlebody of genus $k$ is embedded in the interior of a 4-manifold. An operation of removing the 1-handlebody from the 4-manifold and gluing $\natural_{k}S^{2}\times D^{2}$ back along the resulting boundary component will be called a surgery along a $4$-dimensional $1$-handlebody. The reverse operation will be called a surgery along an embedded copy of $\natural_{k}(S^{2}\times D^{2})$.
@@
(1) A cobordism $R$ from $M$ to $N$ will be called quasi-product if it is orientedly diffeomorphic to the product $I\times M$ possibly after performing surgeries along finitely many embedded $4$-dimensional $1$-handlebodies. If furthermore the homomorphism $H_{2}(M)\to H_{2}(R)$ induced by the inclusion is an isomorphism, then $R$ will be called strongly quasi-product.
(2) A cobordism $P$ from $M$ to $N$ will be called quasi-invertible (resp. strongly quasi-invertible) if it admits a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is a quasi-product cobordism (resp. strongly quasi-product cobordism) from the boundary component $M$ of $P$ to that of $Q$. The cobordism $Q$ will be called a quasi-inverse cobordism (resp. strongly quasi-inverse cobordism) of $P$.
(3) A cobordism $P$ from $M$ to $N$ will be called $H_{2}$-surjective if the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is surjective.
As we will later show, there are numerous examples of quasi-invertible cobordisms. To produce examples using intensively studied objects knotted 2-spheres, let us recall a few terminologies. A surgery on a circle embedded in a 4-manifold is an operation of removing the tubular neighborhood $S^{1}\times D^{3}$ of the circle and gluing $S^{2}\times D^{2}$ back along the resulting boundary component. In general, there are two ways to glue $S^{2}\times D^{2}$ along the boundary component ([33]), but we use a convenient way in each context. Clearly, any surgery on a circle is realized by a surgery along an embedded $4$-dimensional $1$-handlebody of genus one. The surgery on a $2$-sphere embedded in a 4-manifold with trivial normal bundle is an operation of removing the tubular neighborhood $S^{2}\times D^{2}$ of the 2-sphere from the 4-manifold and gluing $S^{1}\times D^{3}$ back along the resulting boundary. We note that the resulting 4-manifold is uniquely determined by the embedded 2-sphere ([47]). An $S^{2}$-link in a 4-manifold is a submanifold of the interior of the 4-manifold whose each connected component is a 2-sphere.
@@
$(1)$ A $4$-manifold is a quasi-product cobordism from $M$ to $M$ if and only if the $4$-manifold is diffeomorphic to a $4$-manifold obtained from $I\times M$ by performing surgeries along finitely many embedded copies of the boundary sums of $S^{2}\times D^{2}$’s.
@@
Now we produce examples of quasi-invertible cobordisms. Clearly, invertible cobordisms are strongly quasi-invertible, and quasi-product cobordisms are quasi-invertible. To produce non-trivial examples, we recall the notion of a strongly slice link.
@@
A (1-dimensional) link $\ell$ in $N$ will be called strongly slice if $\ell$ in $1\times N$ bounds a disjoint union of smoothly and properly embedded disks in $I\times N$. We note that each component of a strongly slice link is null-homologous, and thus its 0-framing is defined as the Seifert framing.
We can generate various examples of quasi-invertible cobordisms from the above trivial examples by the lemma below.
@@
Let $L$ be an $S^{2}$-link in $S^{4}$, and let $V$ be a codimension-00 submanifold of $S^{4}(L)$. Then, for any quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism $P$ from $M$ to $N$, the following hold.
$(1)$ Every codimension-00 submanifold of $P$ containing the boundary component $M$ of $P$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the other boundary component.
$(2)$ $P\#V$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the disjoint union $N\sqcup\partial V$.
$(3)$ If $\partial V$ is non-empty, then $P\natural V$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to $N\#\partial V$, where the boundary sum is taken for the boundary component $N$ of $P$.
$(4)$ Every $4$-manifold obtained by attaching $1$-handles to $P$ along a connected component of $N$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the new boundary component.
$(5)$ Every $4$-manifold obtained by attaching an algebraically canceling pair of $1$- and $2$-handles to $P$ along $N$ is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the new boundary component.
$(6)$ Every $4$-manifold obtained by attaching $2$-handles to $P$ along a strongly slice link $\ell$ in $N$ with 00-framings is a quasi-invertible ((((resp. strongly quasi-invertible)))) cobordism from $M$ to the 00-surgery on $\ell$ in $N$.
@@
The claim (1) is straightforward, and the claims (2) and (3) are straightforward from Lemma 6.4 and the claim (1). For (4), by further attaching a canceling 2-handle to each 1-handle, we obtain a 4-manifold diffeomorphic to $P$. The claim (4) thus follows from (1). For (5), by further attaching the upside down cobordism of the cobordism consisting of the handle pair, we obtain a 4-manifold diffeomorphic to $P$. The claim (5) thus follows from (1). The claim (6) is easily checked as follows. The link $\ell$ in $0\times N$ bounds a disjoint union of smoothly embedded disks in $I\times N$, and the tubular neighborhood of these disks are the 2-handles attached to $P$. Thus the 4-manifold in (6) can be embedded into $P\cup_{\mathrm{id}_{N}}(I\times N)$, and thus the claim (6) follows from (1). ∎
See also Section 8 for examples of codimension-00 submanifolds of $S^{4}(L)$.
@@ 6.2. Algebraic properties
We study algebraic properties of quasi-invertible cobordisms. We begin with quasi-product cobordisms.
@@
Let $\widehat{R}$ be the 4-manifold obtained from $R$ by performing surgeries along finitely many embedded 1-handlebodies $\natural_{k_{i}}S^{1}\times D^{3}$’s so that $\widehat{R}$ is diffeomorphic to $I\times M$. We remove the interior of these 1-handlebodies from $R$ and denote the resulting 4-manifold by $R_{c}$. Let ${j}:M\hookrightarrow{R}$, $j_{c}:M\hookrightarrow R_{c}$, $h:R_{c}\hookrightarrow R$, $\widehat{h}:R_{c}\hookrightarrow\widehat{R}$ and $\widehat{j}:M\hookrightarrow\widehat{R}$ be the inclusions.
By the excision theorem and the Poincaré duality, we see that $H_{3}(\widehat{R},R_{c})\cong\oplus H_{3}(\natural_{k_{i}}(S^{2}\times D^{2}),k_{i}(S^{1}\times S^{2}))\cong 0$. The homology exact sequence for the pair $(\widehat{R},R_{c})$ thus shows that $\widehat{h}_{*}:H_{2}(R_{c})\to H_{2}(\widehat{R})$ is injective. Since the inclusion $\widehat{j}=\widehat{h}\circ j_{c}:M\hookrightarrow\widehat{R}$ is a homotopy equivalence, the injection $\widehat{h}_{*}$ and also ${j_{c}}_{*}:H_{2}(M)\to H_{2}(R_{c})$ are isomorphisms.
@@
If $M$ is a disjoint union of rational homology $3$-spheres, then any quasi-product ((((resp. quasi-invertible)))) cobordism from $M$ to $N$ is strongly quasi-product ((((resp. strongly quasi-invertible)))).
@@
We next discuss quasi-invertible cobordisms. In the rest of this subsection, let $P$ be a quasi-invertible cobordism from $M$ to $N$, and let $Q$ be a quasi-inverse cobordism of $P$. We denote a quasi-product cobordism $P\cup_{N}Q$ by $R$. In the case where $P$ is strongly quasi-invertible, we assume $R$ is strongly quasi-product. Furthermore, let $A^{M}_{P}$ and $B_{P}$ respectively denote the image of the homomorphism $H_{2}(M)\to H_{2}(P)$ and the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$, where these are the homomorphisms induced by the inclusions. Then the following lemma holds.
@@
$(1)$ The homomorphism $H_{1}(M)\to H_{1}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=A^{M}_{P}+B_{P}$ holds.
$(2)$ If $P$ is strongly quasi-invertible, then the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=A^{M}_{P}\oplus B_{P}$ holds.
@@
Finally, we discuss algebraic properties of an attachment of a quasi-invertible cobordisms. Let $X$ be a 4-manifold having a boundary component orientedly diffeomorphic to $M$, and let $X^{\prime}$ denote a 4-manifold $X\cup_{M}P$. We attach $Q$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism. Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$. Let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Then the following properties hold.
@@
$(2)$ $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$.
$(3)$ The restriction of the intersection form of $X^{\prime}$ to ${\iota_{P}}_{*}(B_{P})$ is the zero-map.
@@
$(1)$ By the homology exact sequence for the pair $(R,M)$ and Lemma 6.8, we see $H_{2}(R,M)=0$. The excision theorem shows $H_{2}(X^{\prime\prime},X)\cong H_{2}(R,M)=0$. The homology exact sequence for the pair $(X^{\prime\prime},X)$ thus shows that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is surjective.
$(2)$ The claim (1) and the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$ implies $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+B_{X^{\prime}}$, where $B_{X^{\prime}}$ denotes the kernel of the homomorphism $H_{2}(X^{\prime})\to H_{2}(X^{\prime\prime})$ induced by the inclusion. We note ${\iota_{P}}_{*}(B_{P})\subset B_{X^{\prime}}$. Using the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$ and Lemma 6.10, we can check $\iota_{*}(H_{2}(X))+B_{X^{\prime}}=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$, showing the claim.
@@
$(4)$ By Lemma 6.10 and the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$, the homomorphism $\iota_{*}:H_{1}(X)\to H_{1}(X^{\prime})$ is injective. Thus, due to the homology exact sequence for the pair $(X^{\prime},X)$, the homomorphism $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is surjective if and only if $H_{2}(X^{\prime},X)=0$. Similarly, by Lemma 6.10, $P$ is $H_{2}$-surjective if and only if $H_{2}(P,M)=0$. The claim thus follows from the fact $H_{2}(X^{\prime},X)\cong H_{2}(P,M)$. ∎
The corollary below is straightforward.
@@
We also have the following decomposition.
Lemma 6.13.
If the intersection form of $X$ is non-degenerate ((((on $H_{2}(X)/\mathrm{Tor}$)))), then $H_{2}(X^{\prime})/\mathrm{Tor}=\iota_{*}(H_{2}(X))/\mathrm{Tor}\,\oplus\,{\iota_{P}}_{*}(B_{P})/\mathrm{Tor}$.
Proof.
Since every non-torsion class of $H_{2}(X)$ algebraically intersects with some class of $H_{2}(X)$ under the assumption, the claim easily follows from Lemma 6.11. ∎
@@
Suppose the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is injective. Then, $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is injective, and the decomposition $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(B_{P})$ holds.
@@
The injectivity of $\iota_{*}$ is clear from the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$. Since the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, we easily see $\iota_{*}(H_{2}(X))\cap{\iota_{P}}_{*}(B_{P})=0$. Lemma 6.11 thus shows $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(B_{P})$. ∎
@@
Then the homomorphisms $H_{2}(X)\to H_{2}(X^{\prime})$ and $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusions are injective.
@@
In the case where (i) or (ii) holds, by the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}R$ and Lemma 6.8, we see that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective, and thus the homomorphism $H_{2}(X)\to H_{2}(X^{\prime})$ is also injective. In the case where (iii) holds, we easily see that the condition (ii) also holds, and hence the claim holds. ∎
The property below is useful for studying genus functions.
@@
Then, $H_{2}(P)$ and ${\iota_{P}}_{*}(B_{P})$ are torsion-free.
@@
By the Poincaré duality and the universal coefficient theorem, we see that $H_{2}(N)$ and $H_{2}(M)$ are torsion-free. Since $P$ is strongly quasi-inverible, $H_{2}(R)$ is also torsion-free.
We consider the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$, and we will show that the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ in the sequence is injective. In the case where (i) holds, we may assume that $Q=I\times N$, and hence this homomorphism is injective. In the case where (iii) holds, we see $H_{2}(N)=0$ as shown in the proof of Corollary 6.9, and hence the homomorphism is injective. So, we prove the injectivity in the case where (ii) holds. In this case, we may assume $R$ is diffeomorphic to $I\times M$, and hence the inclusion $M\hookrightarrow R$ is a homotopy equivalence. Due to the sequence $M\subset P\subset R$, this fact implies that the homomorphism $H_{3}(P)\to H_{3}(R)$ induced by the inclusion is surjective. The Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ thus shows that the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective.
Now we are ready to prove the claim. Since $H_{2}(N)$ and $H_{2}(R)$ are torsion-free, and the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective, the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ implies that $H_{2}(P)$ is torsion-free, and hence $B_{p}$ is also torsion-free. By Lemma 6.10, we have $H_{2}(P)=A^{M}_{P}\oplus B_{P}$. The Mayer–Vietoris exact sequence for $X^{\prime}=X\cup_{M}P$ thus shows that the restriction of ${\iota_{P}}_{*}:H_{2}(P)\to H_{2}(X^{\prime})$ to $B_{P}$ is an isomorphism. Therefore, ${\iota_{P}}_{*}(B_{P})$ is torsion-free. ∎

### v5

@@ 6.1. Definitions and examples
Suppose a 4-dimensional 1-handlebody of genus $k$ is embedded in the interior of a 4-manifold. An operation of removing the 1-handlebody from the 4-manifold and gluing $\natural_{k}S^{2}\times D^{2}$ back along the resulting boundary component is called a surgery along a $4$-dimensional 11-handlebody. The reverse operation is called a surgery along an embedded copy of $\natural_{k}(S^{2}\times D^{2})$.
@@
(1) A cobordism $R$ from $M$ to $N$ is called quasi-product if it is orientedly diffeomorphic to the product $I\times M$ possibly after performing surgeries along finitely many embedded $4$-dimensional 11-handlebodies. If furthermore the homomorphism $H_{2}(M)\to H_{2}(R)$ induced by the inclusion is an isomorphism, then $R$ is called strongly quasi-product.
(2) A cobordism $P$ from $M$ to $N$ is called quasi-invertible (resp. strongly quasi-invertible) if it admits a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is a quasi-product cobordism (resp. strongly quasi-product cobordism) from the boundary component $M$ of $P$ to that of $Q$. The cobordism $Q$ is called a quasi-inverse cobordism (resp. strongly quasi-inverse cobordism) of $P$.
(3) A cobordism $P$ from $M$ to $N$ is called $H_{2}$-surjective if the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is surjective.
As we will later show, there are rich examples of quasi-invertible cobordisms. To produce examples, let us recall a few terminologies. A surgery on a circle embedded in a 4-manifold is an operation of removing the tubular neighborhood $S^{1}\times D^{3}$ of the circle and gluing $S^{2}\times D^{2}$ back along the resulting boundary component. In general, there are two ways to glue $S^{2}\times D^{2}$ along the boundary component ([34]), but we use a convenient way in each context. Clearly, any surgery on a circle is realized by a surgery along an embedded $4$-dimensional 11-handlebody of genus one. The surgery on a $2$-sphere embedded in a 4-manifold with trivial normal bundle is an operation of removing the tubular neighborhood $S^{2}\times D^{2}$ of the 2-sphere from the 4-manifold and gluing $S^{1}\times D^{3}$ back along the resulting boundary. We note that the resulting 4-manifold is uniquely determined by the embedded 2-sphere ([49]). An $S^{2}$-link in a 4-manifold is a submanifold of the interior of the 4-manifold whose each connected component is a 2-sphere.
@@
$(1)$ A $4$-manifold is a quasi-product cobordism from $M$ to $M$ if and only if the $4$-manifold is diffeomorphic to a $4$-manifold obtained from $I\times M$ by performing surgeries along finitely many submanifolds, each of which is diffeomorphic to the boundary sum of copies of $S^{2}\times D^{2}$.
@@
Now we produce examples of quasi-invertible cobordisms. Clearly, invertible cobordisms are strongly quasi-invertible, and quasi-product cobordisms are quasi-invertible. To produce varieties of examples, we recall the notion of a strongly slice link.
@@
A (1-dimensional) link $\ell$ in $N$ is called strongly slice if $\ell$ in $1\times N$ bounds a disjoint union of smoothly and properly embedded disks in $I\times N$. We note that each component of a strongly slice link is null-homologous, and thus its 0-framing is defined as the Seifert framing.
In Subsection 8.1, we will produce various examples of codimension-0 submanifolds of $S^{4}(L)$. Using those examples, Lemma 6.4 and the lemma below, we can generate various examples of (strongly) quasi-invertible cobordisms from product cobordisms.
@@
Let $L$ be an $S^{2}$-link in $S^{4}$, and let $V$ be a codimension-0 submanifold of $S^{4}(L)$. Then, for any quasi-invertible ((resp. strongly quasi-invertible)) cobordism $P$ from $M$ to $N$, the following hold.
$(1)$ Every codimension-0 submanifold of $P$ containing the boundary component $M$ of $P$ is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to the other boundary component.
$(2)$ P#VP\#V is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to the disjoint union $N\sqcup\partial V$.
$(3)$ If $\partial V$ is non-empty, then $P\natural V$ is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to $N\#\partial V$, where the boundary sum is taken for the boundary component $N$ of $P$.
$(4)$ Every $4$-manifold obtained by attaching $1$-handles to $P$ along a connected component of $N$ is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to the new boundary component.
$(5)$ Every $4$-manifold obtained by attaching homotopically canceling pairs of $1$- and $2$-handles to $P$ along $N$ is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to the new boundary component.
$(6)$ Every $4$-manifold obtained by attaching $2$-handles to $P$ along a strongly slice link $\ell$ in $N$ with 0-framings is a quasi-invertible ((resp. strongly quasi-invertible)) cobordism from $M$ to the 0-surgery on $\ell$ in $N$.
@@
For the claim (1), attach the quasi-inverse cobordism of $P$ to the exterior of the submanifold. The resulting manifold gives a quasi-inverse cobordism of the submanifold, implying the claim (1). The claims (2) and (3) are straightforward from Lemma 6.4 and the claim (1). For (4), by further attaching a canceling 2-handle to each 1-handle, we obtain a 4-manifold diffeomorphic to $P$. The claim (4) thus follows from (1). For (5), by further attaching the upside down cobordism of the cobordism consisting of the handle pairs, we obtain a 4-manifold diffeomorphic to $P$ due to Lemma 3.4. The claim (5) thus follows from (1). The claim (6) is easily checked as follows. The link $\ell$ in $0\times N$ bounds a disjoint union of smoothly embedded disks in $I\times N$, and the tubular neighborhood of these disks give the 2-handles attached to $P$. Hence, the 4-manifold in (6) can be embedded into $P\cup_{\mathrm{id}_{N}}(I\times N)$, and thus the claim (6) follows from (1). The last claim easily follows from these constructions. ∎
@@ 6.2. Algebraic properties
We begin with quasi-product cobordisms.
@@
Let $\widehat{R}$ be the 4-manifold obtained from $R$ by performing surgeries along finitely many embedded 1-handlebodies $\natural_{k_{i}}S^{1}\times D^{3}$ so that $\widehat{R}$ is diffeomorphic to $I\times M$. We remove the interior of these 1-handlebodies from $R$ and denote the resulting 4-manifold by $R_{c}$. Note that $R=\widehat{R}=R_{c}$ in the case where $R$ is a product cobordism. Let ${j}:M\hookrightarrow{R}$, $j_{c}:M\hookrightarrow R_{c}$, $h:R_{c}\hookrightarrow R$, $\widehat{h}:R_{c}\hookrightarrow\widehat{R}$ and $\widehat{j}:M\hookrightarrow\widehat{R}$ be the inclusions.
[…]
@@
If $M$ is a disjoint union of rational homology $3$-spheres, then any quasi-product ((resp. quasi-invertible)) cobordism from $M$ to $N$ is strongly quasi-product ((resp. strongly quasi-invertible)).
@@
We next discuss quasi-invertible cobordisms. In the rest of this subsection, let $P$ be a quasi-invertible cobordism from $M$ to $N$, and let $Q$ be a quasi-inverse cobordism of $P$. We denote a quasi-product cobordism $P\cup_{N}Q$ by $R$. In the case where $P$ is strongly quasi-invertible, we assume $R$ is strongly quasi-product. Let $I^{M}_{P}$ and $K^{R}_{P}$ respectively denote the image of the homomorphism $H_{2}(M)\to H_{2}(P)$ and the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$, where these are the homomorphisms induced by the inclusions. Then the following lemma holds.
@@
$(1)$ The homomorphism $H_{1}(M)\to H_{1}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=I^{M}_{P}+K^{R}_{P}$ holds.
$(2)$ If $P$ is strongly quasi-invertible, then the homomorphism $H_{2}(M)\to H_{2}(P)$ induced by the inclusion is injective, and the decomposition $H_{2}(P)=I^{M}_{P}\oplus K^{R}_{P}$ holds.
@@
Finally, we discuss algebraic effects of an attachment of a quasi-invertible cobordisms. Let $X$ be a 4-manifold having a boundary component orientedly diffeomorphic to $M$, and let $X^{\prime}$ denote a 4-manifold $X\cup_{M}P$. We attach $Q$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism. In the case where $P$ is strongly quasi-invertible (resp. invertible), we assume $R$ is strongly quasi-product (resp. product). Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$, and let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Then the following properties hold.
@@
$(2)$ $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(K^{R}_{P})$.
$(3)$ The restriction of the intersection form of $X^{\prime}$ to ${\iota_{P}}_{*}(K^{R}_{P})$ is the zero-map.
@@
$(1)$ By the homology exact sequence for the pair $(R,M)$ and Lemma 6.8, we see $H_{2}(R,M)=0$. The excision theorem shows $H_{2}(X^{\prime\prime},X)\cong H_{2}(R,M)=0$. The homology exact sequence for the pair $(X^{\prime\prime},X)$ thus shows the claim.
$(2)$ The claim (1) and the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$ implies $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+K^{X^{\prime\prime}}_{X^{\prime}}$, where $K^{X^{\prime\prime}}_{X^{\prime}}$ denotes the kernel of the homomorphism $H_{2}(X^{\prime})\to H_{2}(X^{\prime\prime})$ induced by the inclusion. We note ${\iota_{P}}_{*}(K^{R}_{P})\subset K^{X^{\prime\prime}}_{X^{\prime}}$. Using the Mayer–Vietoris exact sequence for the decomposition $X^{\prime}=X\cup_{M}P$ and Lemma 6.10, we can check $\iota_{*}(H_{2}(X))+K^{X^{\prime\prime}}_{X^{\prime}}=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(K^{R}_{P})$, showing the claim.
@@
$(4)$ Lemma 6.10 and the Mayer–Vietoris exact sequence for $X^{\prime}=X\cup_{M}P$ imply that $\iota_{*}:H_{1}(X)\to H_{1}(X^{\prime})$ is injective. It thus follows from the homology exact sequence for the pair $(X^{\prime},X)$ that $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is surjective if and only if $H_{2}(X^{\prime},X)=0$. Similarly, by Lemma 6.10, $P$ is $H_{2}$-surjective if and only if $H_{2}(P,M)=0$. The claim thus follows from the fact $H_{2}(X^{\prime},X)\cong H_{2}(P,M)$. ∎
The useful corollary below is straightforward from the above (4).
@@
Suppose the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is injective. Then, $\iota_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is injective, and the decomposition $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))\oplus{\iota_{P}}_{*}(K^{R}_{P})$ holds.
@@
The injectivity is clear from the sequence $X\subset X^{\prime}\subset X^{\prime\prime}$, and the decomposition easily follows from Lemma 6.11. ∎
@@
Then the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ induced by the inclusion is injective.
@@
In the case where (i) or (ii) holds, the Mayer–Vietoris exact sequence for $X^{\prime\prime}=X\cup_{M}R$ implies that the homomorphism $H_{2}(X)\to H_{2}(X^{\prime\prime})$ is injective. In the case where (iii) holds, every class of $H_{2}(X)$ algebraically intersects with some class of $H_{2}(X)$, implying that the condition (ii) also holds. Hence, the claim follows. ∎
The lemma below is useful for applying the results in Section 7.
@@
Then, $H_{2}(P)$ and ${\iota_{P}}_{*}(K^{R}_{P})$ are torsion-free.
@@
The Poincaré duality and the universal coefficient theorem imply that $H_{2}(N)$ and $H_{2}(M)$ are torsion-free. Since $P$ is strongly quasi-inverible, we see that $H_{2}(R)$ is also torsion-free.
We consider the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$, and we will show that the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ in the sequence is injective. In the case where (i) holds, we may assume that $Q=I\times N$, and hence this homomorphism is injective. In the case where (iii) holds, we see $H_{2}(N)=0$ as shown in the proof of Corollary 6.9, and hence the homomorphism is injective. So, we prove the injectivity in the case where (ii) holds. In this case, $R$ is diffeomorphic to $I\times M$, and hence the inclusion $M\hookrightarrow R$ is a homotopy equivalence. Due to the sequence $M\subset P\subset R$, this fact implies that the homomorphism $H_{3}(P)\to H_{3}(R)$ induced by the inclusion is surjective. The Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ thus shows that the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective.
Now we are ready to prove the claim. Since $H_{2}(N)$ and $H_{2}(R)$ are torsion-free, and the homomorphism $H_{2}(N)\to H_{2}(P)\oplus H_{2}(Q)$ is injective, the Mayer–Vietoris exact sequence for $R=P\cup_{N}Q$ implies that $H_{2}(P)$ is torsion-free, and hence $K^{R}_{P}$ is also torsion-free. By Lemma 6.10, we have $H_{2}(P)=I^{M}_{P}\oplus K^{R}_{P}$. It thus follows from the Mayer–Vietoris exact sequence for $X^{\prime}=X\cup_{M}P$ that the restriction ${\iota_{P}}_{*}|_{K^{R}_{P}}:K^{R}_{P}\to{\iota_{P}}_{*}(K^{R}_{P})$ is an isomorphism. Therefore, ${\iota_{P}}_{*}(K^{R}_{P})$ is torsion-free. ∎
In the case where the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ is injective, we still have nice properties.
Proposition 6.16.
Suppose the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ induced by the inclusion is injective. Then, $\iota_{*}:H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime})/\mathrm{Tor}$ is injective, and the decomposition $H_{2}(X^{\prime})/\mathrm{Tor}=\iota_{*}(H_{2}(X))/\mathrm{Tor}\oplus{\iota_{P}}_{*}(K^{R}_{P})/\mathrm{Tor}$ holds.
Proof.
The proof is similar to that of Proposition 6.13. ∎
We give sufficient conditions that the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ is injective.
Proposition 6.17.
Suppose at least one of the following conditions (i)(iii) holds.
  1. (i)
    $P$ is strongly quasi-invertible.
  2. (ii)
    The homomorphism $H_{2}(M)\to H_{2}(X)/\mathrm{Tor}$ induced by the inclusion is the zero-map.
  3. (iii)
    The intersection form of $X$ is non-degenerate.
Then the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ induced by the inclusion is injective.
Proof.
The proof is similar to that of Proposition 6.14. ∎